OpenStudy (anonymous):

When a positive number is multiplied by the sum of twice the number and half the number, the result is the original number. What is the number? Express your answer as a common fraction. When a positive number is multiplied by the sum of twice the number and half the number, the result is the original number. What is the number? Express your answer as a common fraction. @Mathematics

OpenStudy (anonymous):

a*(2a+1/2a)=a

OpenStudy (anonymous):

Let x denote the number we're looking for. From the problem, \[ x * (2x + \frac{x}{2}) = x \] So we have that \[2x^2 + \frac{x^2}{2} = x \rightarrow \frac{4x^2}{2}+\frac{x^2}{2} = \frac{5x^2}{2} = x\] since we're given that x must be positive, we know it can't be zero. Therefore, it's safe to divide by x, and we find that \[ \frac{5x}{2} = 1 \rightarrow x = \frac{2}{5} \] Note that, if we didn't have the restriction that x must be positive, zero would work just as well.

OpenStudy (kainui):

Simplify it by showing that a(5/2a)=a 5/2a=1 a=2/5